# PSformer: Parameter-efficient Transformer with Segment Attention for Time Series Forecasting

### v1

@@ Abstract
Time series forecasting remains a critical challenge across various domains, often complicated by high-dimensional data and long-term dependencies. This paper presents a novel transformer architecture for time series forecasting, incorporating two key innovations: parameter sharing (PS) and Spatial-Temporal Segment Attention (SegAtt). We also define the time series segment as the concatenation of sequence patches from the same positions across different variables. The proposed model, PSformer, reduces the number of training parameters through the parameter sharing mechanism, thereby improving model efficiency and scalability. The introduction of SegAtt could enhance the capability of capturing local spatio-temporal dependencies by computing attention over the segments, and improve global representation by integrating information across segments. The combination of parameter sharing and SegAtt significantly improves the forecasting performance. Extensive experiments on benchmark datasets demonstrate that PSformer outperforms popular baselines and other transformer-based approaches in terms of accuracy and scalability, establishing itself as an accurate and scalable tool for time series forecasting.
footnotetext: We will open-source the code upon acceptance.footnotetext: * Equal contribution.footnotetext: † Corresponding authors.
@@ 1 Introduction
Time series forecasting is an important learning task with significant application values in a wide range of domains, including the weather prediction (Ren et al., 2021; Chen et al., 2023a), traffic flow (Tedjopurnomo et al., 2020; Khan et al., 2023), energy consumption (Liu et al., 2020; Nti et al., 2020), anomaly detection (Zamanzadeh Darban et al., 2022) and the financial analysis (Nazareth & Ramana Reddy, 2023), etc. With the advancement of artificial intelligence techniques, significant efforts have been devoted to developing innovative models that continue to improve the prediction performance (Liang et al., 2024; Wang et al., 2024). In particular, the transformer-based model family has recently attracted more attention for its proved success in nature language processing (OpenAI et al., 2024) and computer vision (Liu et al., 2021; Dosovitskiy et al., 2021). Moreover, pre-trained large models based on the transformer architecture have shown advantages in time series forecasting(Liu et al., 2024a; Jin et al., 2024; Chang et al., 2023; Woo et al., 2024), demonstrating that increasing the amount of parameters in transformer models and the volume of training data can effectively enhance the model capability.
@@
On the other side, many simple linear models (Zeng et al., 2023; Li et al., 2023) also shown competitive performance compared to the more complex designs of transformer-based models. One possible key reason for their success in time series forecasting is that they have low model complexities and are less likely to overfit on noisy or irrelevant signals. As a result, even with limited data, these models can effectively capture robust information representations. To overcome the limitations of modeling long-term dependencies and capturing complex temporal relationships, PatchTST (Nie et al., 2023) process temporal information by combining patching techniques to extract local semantic information, leading to superior performance. However, it applies channel-independent designs and leaves the significant potential for improvement in effectively modeling across variables. Moreover, the unique challenges of modeling multivariate time series data, where the temporal and spatial dimensions differ significantly from other data types, present many unexplored opportunities. While past research (donghao & wang xue, 2024; Zhang & Yan, 2023; Ilbert et al., 2024) has largely treated these dimensions separately, the question of whether systematically mixing temporal and spatial information can further enhance model performance remains an open area for future investigation.
@@
To reduce the model complexity in deep learning, parameter sharing (PS) is a crucial technique that can significantly reduce the amount model parameters and enhance computational efficiency. In CNNs, convolutional filters share weights across spatial locations, capturing local features with fewer parameters. Similarly, LSTM networks share weight matrices across time steps to manage memory and control information flow. By studying the sharing of attention weights, (Xiao et al., 2019; Kitaev et al., 2020) improved performance in the fields of natural language processing and computer vision. ALBERT (Lan et al., 2020) further extends parameter sharing in natural language processing by sharing weights across transformer layers, reducing parameter redundancy while maintaining performance. In multi-task learning, the Task Adaptive Parameter Sharing (TAPS) approach (Wallingford et al., 2022) selectively fine-tunes task-specific layers while maximizing parameter sharing across tasks, achieving efficient learning with minimal task-specific modifications. Those studies demonstrate that parameter sharing has the potential for model size reduction, generalization ability enhancement and mitigating the over-fitting risks across various tasks. Besides, employing advanced optimization technique is also essential in reducing the overfitting issue. (Ilbert et al., 2024) proposed a simple yet effective transformer-based network structure and equipped it with the SAM (Sharpness-Aware Minimization) optimization mechanism to achieve competitive performance compared with large models (Foret et al., 2021).
@@
In this work, we explore innovative designs of transformer-based model for time series forecasting by incorporating the characteristics of time series tasks and the concept of parameter sharing. Unlike MOIRAI (Woo et al., 2024) that “flattens” multivariate time series by treating all variables as a single sequence, or SAMformer (Ilbert et al., 2024) that applies attention to the channel dimension to capture spatial dependencies, we propose a different and novel approach. We construct a transformer encoder with two-stage segment attention structure, where each network layer consists of a parameter-sharing block. This parameter-sharing block is composed of three fully connected layers with residual connections, which keeps the overall number of parameters in the model very small while enabling effective information sharing across different parts of the model. For segment attention, we use patching to divide the variables into different patches, then identify the patches at the same position across different variables and merge them into a segment. As a result, each segment represents the spatial extension of a single-variable patch. In this way, we decompose the multivariate time series into multiple segments. Attention is applied within each segment to enhance the extraction of local spatial-temporal relationships, while information fusion across segments is performed to improve the overall predictive performance. Additionally, by incorporating the SAM optimization method, we further reduce over-fitting while maintaining efficient training. We conduct extensive experiments on long-term time series forecasting datasets to verify the effectiveness of this segment attention structure with parameter sharing. Our model remains competitive when compared to previous state-of-the-art models, achieving the best performance on 6 out of 8 mainstream long-term time series forecasting tasks. The contributions are summarized as follows:
@@ 2.1 Temporal modeling in time series forecasting
On the other hand, some pre-trained large models have been applied to time series forecasting tasks (Das et al., 2023; Liu et al., 2024a; Gao et al., 2024; Liu et al., 2024c; Zhou et al., 2023; Jin et al., 2024). For example, Moment uses the patching method and mask pre-training to build a pre-trained model for time series, while GPT4TS (Zhou et al., 2023) also adopts the patching method and uses BERT as the backbone. The increase in model parameters has provided them with greater expressive power but also increased the difficulty of training.
@@ 2.2 Variate modeling in time series forecasting
Figure 1: Data transformation in PSformer. Multivariate time series data first undergo patching and cross-channel merging before being fed into the PSformer Encoder. Then after inverse transformation and linear mapping for the prediction length, the final time series results are obtained.
@@ 3.2 Model Structure
The PSformer model we constructed is depicted in Figure 2. The key components of the model include the PSformer Encoder, segment attention, and PS Block. The PSformer Encoder serves as the backbone of the model and contains both the segment attention and the PS Block. The PS Block provides the parameters for all layers within the Encoder, utilizing the parameter sharing technique.
Figure 2: PSformer network structure. The PSformer includes the PS Block for parameter sharing within the PS Encoder structure, as well as the two-stage segment attention.
Forward Process The univariate time series of length $L$ for the $i$-th variable, starting at time index $1$, is denoted as $x_{1}^{(i)}=(x_{1}^{(i)},...,x_{L}^{(i)})$, where $i=1,...,M$. Then the input $(x_{1},...,x_{L})$ with $M$ dimensions is presented as $x_{1}\in\mathbb{R}^{M\times L}$, and $x_{1}$ is used as the input to the transformer network structure. Similar to other time series forecasting methods, we use a RevIN layer Kim et al. (2022), which is added at both the input and output of the model.
Spatial-Temporal Segment Attention
We introduce spatial-temporal segment attention (SegAtt), which merges patches from different channels at the same position into a segment and establishes spatial-temporal relationships across different segments. Specifically, the input time series $X\in\mathbb{R}^{M\times L}$ is first divided into patches, where $L=P\times N$, and then transformed into $X\in\mathbb{R}^{(M\times P)\times N}$. By merging the $M$ and $P$ dimensions, the input becomes $X\in\mathbb{R}^{C\times N}$, where $C=M\times P$, facilitating the subsequent cross-channel information fusion.
In this transformed space, identical $\bm{Q}\in\mathbb{R}^{C\times N}$, $\bm{K}\in\mathbb{R}^{C\times N}$, and $\bm{V}\in\mathbb{R}^{C\times N}$ matrices are generated by applying a shared block’s non-linear projection (applied by Parameter Shared Block), with weights $\bm{W}^{Q}\in\mathbb{R}^{N\times N}$ and $\bm{W}^{K}\in\mathbb{R}^{N\times N}$ used to project the input $X^{in}\in\mathbb{R}^{C\times N}$. The $\bm{Q}$ and $\bm{K}$ matrices are then multiplied using a dot-product operation to form the attention matrix $\bm{Q}K^{T}\in\mathbb{R}^{C\times C}$, which captures relationships between different spatial-temporal segments (in the $C$ dimension) and is used to act on the $\bm{V}$ matrix. While the computation of $\bm{Q}$, $\bm{K}$ and $\bm{V}$ involves nolinear transformations of the input $X^{in}$ across segments in the $N$ dimension (corresponding to the $d_{m}odel$ dimension of attention in the NLP domain), the scaled dot-product attention primarily applies attention across the $C$ dimension, allowing the model to focus on dependencies between spatial-temporal segments across channels and time.
@@
$\bm{X}^{out}=\text{GeLU}(\bm{X}^{in}\bm{W}^{(1)})\bm{W}^{(2)}+\bm{X}^{in},\vspace{0.5ex}$ (1)
which follows a similar structure as the feed-forward network (FFN) with residual connections. This intermediate output $\bm{X}^{out}$ is then used as the input for the third transformation, yielding:
$\bm{X}^{out}=\bm{X}^{in}\bm{W}^{(3)}.$
@@
and we denote PS Block output as $\bm{X}^{out}=\bm{X}^{in}\bm{W}^{S}$, where $W^{S}\in\mathbb{R}^{N\times N}$ and $\bm{X}^{out}\in\mathbb{R}^{C\times N}$. The structure of the PS Block allows it to perform nonlinear transformations while preserving a linear transformation path. Although the three layers within the PS Block have different parameters, the entire PS Block is reused across different positions in the PSformer Encoder, ensuring that the same block parameters $\bm{W}^{S}$ are shared across those positions, as illustrated in Figure 2. Specifically, PS Block share parameters across three parts of each PSformer Encoder, which including two segment attention layer and the final PS Block. In the segment attention layer, the PS block outputs are used as the $Q$, $K$, and $V$ matrices to construct the attention mechanism. This parameter-sharing strategy reduces the overall number of parameters while maintaining the network’s expressive capacity.
@@
PSformer Encoder In the PSformer Encoder, as Figure 2, each layer shares the same parameters $\bm{W}^{S}$ of PS Block. For the input $\bm{X}^{in}$, the transformation in the PSformer Encoder can be expressed as follows:
@@
SegAtt stage one is represented as: $\bm{Q}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, $\bm{K}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, $\bm{V}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, Therefore, $\bm{Q}^{(1)}$, $\bm{K}^{(1)}$, $\bm{V}^{(1)}$ $\in\mathbb{R}^{C\times N}$. Dot-product attention operation
$\bm{O}^{(1)}=\text{Attention}^{(1)}(\bm{Q}^{(1)},\bm{K}^{(1)},\bm{V}^{(1)})=\text{Softmax}(\frac{\bm{Q}^{(1)}\bm{K}^{(1)}}{\sqrt{d_{k}}}){V}^{(1)}.$
then with ReLU activation:$\bm{O}^{(1)}_{act}=ReLU(\bm{O}^{(1)})$
@@
SegAtt stage two is represented as: $\bm{Q}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S}$, $\bm{K}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S}$, $\bm{V}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S},$ Therefore, $\bm{Q}^{(2)}$, $\bm{K}^{(2)}$, $\bm{V}^{(2)}$ $\in\mathbb{R}^{C\times N}.$ Dot-product attention operation
$\bm{O}^{(2)}=\text{Attention}^{(2)}(\bm{Q}^{(2)},\bm{K}^{(2)},\bm{V}^{(2)})=\text{Softmax}(\frac{\bm{Q}^{(2)}\bm{K}^{(2)}}{\sqrt{d_{k}}})\bm{V}^{(2)}.$
@@
The two-stage SegAtt mechanism can be viewed as analogous to an FFN layer, where the MLP is replaced with attention operations. Additionally, residual connections are introduced between the input and output, and the result is then fed into the final PS Block.
The transformation in the final PS Block is represented as $\bm{O}^{out}=\bm{O}^{in}\bm{W}^{S}.$ Therefore, the entire encoder can be expressed as
$\bm{X}^{out}=(\text{Attention}^{(2)}(\text{ReLU}(\text{Attention}^{(1)}(\bm{X}^{in})))+\bm{X}^{in})\bm{W}^{S},$
@@
with $\bm{X}^{out}\in\mathbb{R}^{C\times N}$. Then we apply a dimensionality transformation to obtain $\bm{X}^{out}\in\mathbb{R}^{M\times L}$, since $C=M\times P$ and $L=P\times N$.
@@
After passing through $n$ layers of the PSformer Encoder, the final output is $\bm{X}^{pred}=\bm{X}^{out}\bm{W}^{F}$, where $\bm{X}^{pred}\in\mathbb{R}^{M\times F}$, and $\bm{W}^{F}\in\mathbb{R}^{L\times F}$ is a linear mapping, where $F$ is the prediction length. The $\bm{X}^{pred}$ is the final output of the PSformer model.
4 Experiment
Datasets In this paper, we focus on the long-term time series forecasting. We follow the time series forecasting work in Ilbert et al. (2024) and use 8 mainstream datasets to evaluate the performance of our proposed PSformer model. As shown in Table 1, these datasets include 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), as well as Weather, Traffic, Electricity, and Exchange. These datasets have been used as benchmark evaluations in many previous time series forecasting studies.
@@ 4 Experiment
Baselines We select state-of-the-art (SOTA) models in the field of long-term time series forecasting, including not only Transformer-based models but also large models and other SOTA models. Specifically, baselines include (1) Transformer-based model: SAMformer (Ilbert et al., 2024), iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), (2) Pretrained Large model: MOMENT (Goswami et al., 2024), GPT4TS (Zhou et al., 2023), (3) TCN-based model: ModernTCN (donghao & wang xue, 2024) and (4) MLP-based methods: TSMixer (Chen et al., 2023b), RLinear (Li et al., 2023).
Experimental Settings The input time series length $T$ is set to 512, and four different prediction lengths $H\in\{96,192,336,720\}$ are used. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE). We train our constructed models using the SAM optimization technique as in Ilbert et al. (2024). We collect baseline results of SAMformer, TSMixer, FEDformer and Autoformer from Ilbert et al. (2024), in which SAMformer and TSMixer are also trained with SAM, and the results of iTransformer, RLinear and PatchTST from Liu et al. (2024b). We test ModernTCN with the fixed input length of 512 following donghao & wang xue (2024). For the pre-trained large models MOMENT and GPT4TS, the results are collected from (Goswami et al., 2024). More details about the baseline settings can be found in Section A.1.
@@
Table 2: Long-term forecasting task. All the results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better performance. See Table 27 in Appendix for the full results with more baselines.

### v2

@@ Abstract
Time series forecasting remains a critical challenge across various domains, often complicated by high-dimensional data and long-term dependencies. This paper presents a novel transformer architecture for time series forecasting, incorporating two key innovations: parameter sharing (PS) and Spatial-Temporal Segment Attention (SegAtt). The proposed model, PSformer, reduces the number of training parameters through the parameter sharing mechanism, thereby improving model efficiency and scalability. We define the time series segment as the concatenation of sequence patches from the same positions across different variables. The introduction of SegAtt could enhance the capability of capturing local spatio-temporal dependencies by computing attention over the segments, and improve global representation by integrating information across segments. Consequently, the combination of parameter sharing and SegAtt significantly improves the forecasting performance. Extensive experiments on benchmark datasets demonstrate that PSformer outperforms popular baselines and other transformer-based approaches in terms of accuracy and scalability, establishing itself as an accurate and scalable tool for time series forecasting.
Machine Learning, ICML
@@ 1 Introduction
Time series forecasting is an important learning task with significant application values in a wide range of domains, including the weather prediction (Ren et al., 2021; Chen et al., 2023a), traffic flow (Tedjopurnomo et al., 2020; Khan et al., 2023), energy consumption (Liu et al., 2020; Nti et al., 2020), anomaly detection (Zamanzadeh Darban et al., 2022) and the financial analysis (Nazareth & Ramana Reddy, 2023), etc. With the advancement of artificial intelligence techniques, significant efforts have been devoted to developing innovative models that continue to improve the prediction performance (Liang et al., 2024; Wang et al., 2024b). In particular, the transformer-based model family has recently attracted more attention for its proved success in nature language processing (OpenAI et al., 2024) and computer vision (Liu et al., 2021; Dosovitskiy et al., 2021). Moreover, pre-trained large models based on the transformer architecture have shown advantages in time series forecasting(Liu et al., 2024a; Jin et al., 2024; Chang et al., 2023; Woo et al., 2024), demonstrating that increasing the amount of parameters in transformer models and the volume of training data can effectively enhance the model capability.
@@
On the other side, many simple linear models (Zeng et al., 2023; Li et al., 2023) also show competitive performance compared to the more complex designs of transformer-based models. One possible key reason for their success in time series forecasting is that they have low model complexities and are less likely to overfit on noisy or irrelevant signals. As a result, even with limited data, these models can effectively capture robust information representations. To overcome the limitations of modeling long-term dependencies and capturing complex temporal relationships, PatchTST (Nie et al., 2023) process temporal information by combining patching techniques to extract local semantic information, leading to superior performance. However, it applies channel-independent designs and leaves the significant potential for improvement in effectively modeling across variables. Moreover, the unique challenges of modeling multivariate time series data, where the temporal and spatial dimensions differ significantly from other data types, present many unexplored opportunities. While past research (donghao & wang xue, 2024; Zhang & Yan, 2023; Ilbert et al., 2024) has largely treated these dimensions separately, the question of whether systematically mixing temporal and spatial information can further enhance model performance remains an open area for future investigation.
@@
To reduce the model complexity in deep learning, parameter sharing (PS) is a crucial technique that can significantly reduce the amount model parameters and enhance computational efficiency. In CNNs, convolutional filters share weights across spatial locations, capturing local features with fewer parameters. Similarly, LSTM networks share weight matrices across time steps to manage memory and control information flow. By studying the sharing of attention weights, (Xiao et al., 2019; Kitaev et al., 2020) improved performance in the fields of natural language processing and computer vision. ALBERT (Lan et al., 2020) further extends parameter sharing in natural language processing by sharing weights across transformer layers, reducing parameter redundancy while maintaining performance. In multi-task learning, the Task Adaptive Parameter Sharing (TAPS) approach (Wallingford et al., 2022) selectively fine-tunes task-specific layers while maximizing parameter sharing across tasks, achieving efficient learning with minimal task-specific modifications. Those studies demonstrate that parameter sharing has the potential for model size reduction, generalization ability enhancement and mitigating the over-fitting risks across various tasks. Besides, employing advanced optimization technique is also essential in reducing the overfitting issue. (Ilbert et al., 2024) proposed a simple yet effective transformer-based network and equipped it with the SAM (Sharpness-Aware Minimization) optimization mechanism to achieve competitive performance compared with large models (Foret et al., 2021).
@@
In this work, we explore innovative designs of the transformer-based model for time series forecasting by incorporating the spatiotemporal dependencies characteristic of time series tasks and a new parameter sharing model architecture. Some past works have proposed methods to capture spatiotemporal dependency information, such as MOIRAI (Woo et al., 2024), which “flattens” multivariate time series by treating all variables as a single sequence, and SAMformer (Ilbert et al., 2024), which applies attention to the channel dimension to capture spatial dependencies. Unlike previous methods, we construct a transformer-based model called PSformer, which incorporates a two-stage segment attention structure, each segment attention consists of a parameter-sharing design. This parameter-sharing design consists of three fully connected layers with residual connections, which keeps the overall number of parameters in both the two-stage segment attention and final fusion stage of the PSformer efficient, while also facilitating information sharing between different modules of the model. For segment attention, we use patching to divide the variables into different patches, then identify the patches at the same position across different variables and merge them into a segment. As a result, each segment represents the spatial extension of a single-variable patch. In this way, we decompose the multivariate time series into multiple segments. Attention is applied within each segment to enhance the extraction of local spatial-temporal relationships, while information fusion across segments is performed to improve the overall predictive performance. Additionally, by incorporating the SAM optimization method, we further reduce over-fitting while maintaining efficient training. We conduct extensive experiments on long-term time series forecasting datasets to verify the effectiveness of this segment attention structure with parameter sharing. Our model demonstrates strong competency in comparison to previous state-of-the-art models, achieving the best performance on 7 out of 8 mainstream long-term time series forecasting tasks. The contributions are summarized as follows:
@@
Figure 1: PSformer Dataflow Pipline. Multivariate time series data first undergo patching and cross-channel merging before being fed into the PSformer Encoder. During the cross-channel patching stage, patches from the same position across different time series variables are concatenated to form a mixed local multivariate sequence, where the different mixed time series maintain a strict chronological order, capturing global spatiotemporal dynamics. Finally, inverse transformation and linear mapping are applied to generate the prediction.
@@ 2.1 Temporal modeling in time series forecasting
On the other hand, some pre-trained large models have been applied to time series forecasting tasks (Das et al., 2023; Liu et al., 2024a; Gao et al., 2024; Liu et al., 2024c; Zhou et al., 2023; Jin et al., 2024). For example, MOMENT (Goswami et al., 2024) uses the patching method and mask pre-training to build a pre-trained model for time series, while GPT4TS (Zhou et al., 2023) also adopts the patching method and uses GPT2 as the backbone. The increase in model parameters has provided them with greater expressive power but also increased the difficulty of training.
@@ 3.1 Problem Formulation
Figure 2: PSformer Network Structure. The PSformer processes input through a two-stage Segment Attention connected via residual structures. In the final fusion stage, it passes through a PS Block to generate the output. Notably, the PS Block used here is identical to those in the two-stage Segment Attention. Within the Segment Attention modules, the outputs of the PS Block are leveraged as Q, K, V matrices to compute cross-channel dot-product attention over temporal sequences. This attention matrix captures local spatiotemporal interactions across channels. The right portion of the figure illustrates this cross-channel transformation mechanism.
@@ 3.2 Model Structure
The constructed PSformer model is depicted in Figure 2, where the PSformer Encoder serves as the backbone of the model and the key components of the Encoder include the segment attention and PS Block. The PS Block provides the parameters for all layers within the Encoder by utilizing the parameter sharing technique.
Forward Process The univariate time series of length $L$ for the $i$-th variable, starting at time index $1$, is denoted as $x_{1}^{(i)}=(x_{1}^{(i)},...,x_{L}^{(i)})$, where $i=1,...,M$. Then the input $(x_{1},...,x_{L})$ with $M$ dimensions is presented as $x_{1}\in\mathbb{R}^{M\times L}$, and $x_{1}$ is used as the input to the transformer network structure. Similar to other time series forecasting methods, we use a RevIN layer (Kim et al., 2022), which is added at both the input and output of the model.
Spatial-Temporal Segment Attention We introduce spatial-temporal segment attention (SegAtt), which merges patches from different channels at the same position into a segment and establishes spatial-temporal relationships across different segments. Specifically, the input time series $X\in\mathbb{R}^{M\times L}$ is first divided into patches, where $L=P\times N$, and then transformed into $X\in\mathbb{R}^{(M\times P)\times N}$. By merging the $M$ and $P$ dimensions, the input becomes $X\in\mathbb{R}^{C\times N}$, where $C=M\times P$, facilitating the subsequent cross-channel information fusion.
In this transformed space, identical $\bm{Q}\in\mathbb{R}^{C\times N}$, $\bm{K}\in\mathbb{R}^{C\times N}$, and $\bm{V}\in\mathbb{R}^{C\times N}$ matrices are generated by applying a shared block’s non-linear projection (applied by Parameter Shared Block), with weights $\bm{W}^{Q}\in\mathbb{R}^{N\times N}$ and $\bm{W}^{K}\in\mathbb{R}^{N\times N}$ used to project the input $X^{in}\in\mathbb{R}^{C\times N}$. The $\bm{Q}$ and $\bm{K}$ matrices are then multiplied using a dot-product operation to form the attention matrix $\bm{Q}K^{T}\in\mathbb{R}^{C\times C}$, which captures relationships between different spatial-temporal segments (in the $C$ dimension) and is used to act on the $\bm{V}$ matrix. While the computation of $\bm{Q}$, $\bm{K}$ and $\bm{V}$ involves non-linear transformations of the input $X^{in}$ across segments in the $N$ dimension (corresponding to the $d_{model}$ dimension of attention in the NLP domain), the scaled dot-product attention primarily applies attention across the $C$ dimension, allowing the model to focus on dependencies between spatial-temporal segments across channels and time.
@@
$\bm{X}^{out}=\text{GeLU}(\bm{X}^{in}\bm{W}^{(1)})\bm{W}^{(2)}+\bm{X}^{in},$ (1)
which follows a similar structure as the feed-forward network (FFN) with residual connections. This intermediate output $\bm{X}^{out}$ is then used as the input for the third transformation, yielding
$\bm{X}^{out}=\bm{X}^{in}\bm{W}^{(3)}.$ (2)
@@
And we denote PS Block output as $\bm{X}^{out}=\bm{X}^{in}\bm{W}^{S}$, where $W^{S}\in\mathbb{R}^{N\times N}$ and $\bm{X}^{out}\in\mathbb{R}^{C\times N}$. The structure of the PS Block allows it to perform nonlinear transformations while preserving a linear transformation path. Although the three layers within the PS Block have different parameters, the entire PS Block is reused across different positions in the PSformer Encoder, ensuring that the same block parameters $\bm{W}^{S}$ are shared across those positions, as illustrated in Figure 2. Specifically, PS Block share parameters across three parts of each PSformer Encoder, which including two segment attention layer and the final PS Block. In the segment attention layer, the PS block outputs are used as the $Q$, $K$, and $V$ matrices to construct the attention mechanism. This parameter-sharing strategy reduces the overall number of parameters while maintaining the network’s expression capacity.
@@
PSformer Encoder In the PSformer Encoder, as shown in Figure 2, each layer shares the same parameters $\bm{W}^{S}$ of PS Block. For the input $\bm{X}^{in}$, the transformation in the PSformer Encoder can be expressed as follows:
@@
SegAtt stage one is represented as: $\bm{Q}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, $\bm{K}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, $\bm{V}^{(1)}=\bm{X}^{in}\bm{W}^{S}$, Therefore, we have $\bm{Q}^{(1)}$, $\bm{K}^{(1)}$, $\bm{V}^{(1)}$ $\in\mathbb{R}^{C\times N}$. The dot-product attention operation can be formulated by
$\bm{O}^{(1)}$ $=\text{Attention}^{(1)}(\bm{Q}^{(1)},\bm{K}^{(1)},\bm{V}^{(1)})$ (4)
$=\text{Softmax}(\frac{\bm{Q}^{(1)}\bm{K}^{(1)}}{\sqrt{d_{k}}}){V}^{(1)},$ (5)
which is followed by the ReLU activation:$\bm{O}^{(1)}_{act}=ReLU(\bm{O}^{(1)})$.
@@
SegAtt stage two is represented as: $\bm{Q}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S}$, $\bm{K}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S}$, $\bm{V}^{(2)}=\bm{O}^{(1)}_{act}\bm{W}^{S}.$ Similary, we have $\bm{Q}^{(2)}$, $\bm{K}^{(2)}$, $\bm{V}^{(2)}$ $\in\mathbb{R}^{C\times N},$ and the dot-product attention operation
$\bm{O}^{(2)}$ $=\text{Attention}^{(2)}(\bm{Q}^{(2)},\bm{K}^{(2)},\bm{V}^{(2)})$ (6)
$=\text{Softmax}(\frac{\bm{Q}^{(2)}\bm{K}^{(2)}}{\sqrt{d_{k}}})\bm{V}^{(2)}.$ (7)
@@
The two-stage SegAtt mechanism can be viewed as analogous to an FFN layer, where the MLP is replaced with attention operations. Additionally, residual connections are introduced between the input and output, and the result is then fed into the final PS Block. Since the transformation in the final PS Block is represented as $\bm{O}^{out}=\bm{O}^{in}\bm{W}^{S},$ the entire encoder can be expressed as
$\small\bm{X}^{out}=(\text{Attention}^{(2)}(\text{ReLU}(\text{Attention}^{(1)}(\bm{X}^{in})))+\bm{X}^{in})\bm{W}^{S},$ (8)
@@
with $\bm{X}^{out}\in\mathbb{R}^{C\times N}$. Finally, since $C=M\times P$ and $L=P\times N$, we apply a dimensionality transformation to obtain $\bm{X}^{out}\in\mathbb{R}^{M\times L}$.
@@
After passing through $n$ layers of the PSformer Encoder, the final output is $\bm{X}^{pred}=\bm{X}^{out}\bm{W}^{F}$, where $\bm{X}^{pred}\in\mathbb{R}^{M\times F}$, $\bm{W}^{F}\in\mathbb{R}^{L\times F}$ is a linear mapping and $F$ is the prediction length. The $\bm{X}^{pred}$ is the final output of the PSformer model. It is worth noting that the PSformer structure does not use positional encoding, as the segment attention mixes local spatio-temporal information, for which we provide a more detailed discussion in Section A.6.
@@
Table 2: Long-term forecasting task. All the results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better performance. See Table B.2 in Appendix for the full results with more baselines.